While exploring an abandoned cathedral, you find an old pipe organ. Only one pipe still works, but most of the pipe's harmonic frequencies are absorbed before you can hear them due to overgrowth and debris. The harmonic frequencies you hear are 138 Hz , 230 Hz , and 322 Hz .

a. Is the pipe closed on one end or is it open on both ends?

Closed on one end.
Open on both ends.
This choice is not correct.

b. How long is the pipe?
Answer #1

a. it is a closed pipe which is closed at one end pipe with fundamental frequency..46 Hz

so the frequencies are in the ratio of 46 , 3(46) = 138, 5 (46)= 230 Hz and 7 (46)=322 Hz..respectively..

b. fundamental frequency f = v/4L

         L = v/4f = 342/(4*46)

           = 1.86m